OpenStudy (anonymous):

will fan and give medal if you help me with my mastery test for positive and negative exponents

OpenStudy (gorv):

where is the Q??

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

1/8^2

OpenStudy (anonymous):

just help me understand how to get the answer lol

OpenStudy (gorv):

what we have to do ??

OpenStudy (anonymous):

it just says find an equivalent exponential expression without denominators. keep the same base

OpenStudy (gorv):

so we need to move 8 to numerator

OpenStudy (gorv):

if we move something from denominator numerator thn its power multiplied by minus sign

OpenStudy (anonymous):

this is easy i can help you !!!

OpenStudy (anonymous):

i dont know you tell me i dont understand it

OpenStudy (gorv):

like 1/x

OpenStudy (gorv):

if we move it up

OpenStudy (gorv):

=x^(-1)

OpenStudy (gorv):

1/x^4 =x^(-4)

OpenStudy (anonymous):

well it wouldnt be 1/64 cause ive tried multiple times for other problems and okay @kissmebaby!!34

OpenStudy (gorv):

power was 4 multiplied by minus sign =-4

OpenStudy (anonymous):

@kissmebaby!!!34

OpenStudy (gorv):

we have to keep the base same

OpenStudy (anonymous):

so its -4?

OpenStudy (anonymous):

over 1

OpenStudy (gorv):

we have \[\frac{ 1 }{ x^{19} }\]

OpenStudy (skullpatrol):

$$\Huge \frac{1}{a^n}=a^{-n}$$

OpenStudy (gorv):

now if we move it numerator

OpenStudy (gorv):

\[\frac{ 1 }{ x^{19} }=x^{-19}\]

OpenStudy (anonymous):

8x8 is 64 what would the exponent for it be?

OpenStudy (skullpatrol):

2

OpenStudy (anonymous):

so i just but 64^2 ?

OpenStudy (skullpatrol):

$$\Huge 8 \cdot 8 = 8^2 = 64$$

OpenStudy (anonymous):

so just 64?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it says

OpenStudy (anonymous):

find an equivalent exponential expression without denominators. keep the same base

OpenStudy (skullpatrol):

do you know what a denominator is?

OpenStudy (anonymous):

yeah its the number on the bottom

OpenStudy (skullpatrol):

$$\Huge \frac{1}{a^n}$$

OpenStudy (gorv):

here base=8

OpenStudy (skullpatrol):

what is the denominator?

OpenStudy (gorv):

we have to keep the base same

OpenStudy (anonymous):

*facepalm , so i put 8/64^2

OpenStudy (gorv):

so keep it as 8

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then...

OpenStudy (gorv):

what ie the power of 8 ??

OpenStudy (anonymous):

umm 64

OpenStudy (skullpatrol):

1/8^2 = 1/(8^2)

OpenStudy (anonymous):

if your saying it as 8x8

OpenStudy (anonymous):

and @skullpatrol that is the question

OpenStudy (anonymous):

im confused 0.0

OpenStudy (gorv):